Site 1124
Hole 1124A
The vessel proceeded directly to the GPS coordinates of proposed site SWPAC-9B where the positioning beacon was deployed and the hydrophones and thrusters were lowered. The APC/XCB BHA was assembled using a 9-7/8-in PDC bit and deployed. Hole 1124A was spudded with the APC at 0610 hr on 26 September. The core barrel was full and required a second attempt to recover a mudline.

Hole 1124B
The bit was positioned at 3962.5 mbsl, which was 5 m higher than the bit position of Hole 1124A. The initial attempt at a mud-line core on this hole was thwarted when the core barrel was recovered with a shattered liner and no core. Another core barrel was deployed and Hole 1124B was spudded with the APC at 0507 hr on 26 September. At 0610 hr the second piston core in Hole 1124B was shot with the bit at 3972.0 mbsl (5.4 mbsf). Because of an incomplete stroke, the drill string pressure had to be manually bled off. The core barrel could not be retrieved with 10 kilopounds (kips) of overpull being applied. Various attempts at freeing the stuck core barrel proved unsuccessful. After the sinker bar assembly was recovered with the wireline the top drive was racked and the BHA with the stuck APC barrel was brought back to the surface. At 1745 hr on 26 September, the bit was at the rotary table. The APC corer was found in a fully stroked out position and the bottom section of the inner barrel was extended below the bit and bent ~1 m below the BHA, which prevented the core barrel being retrieved through the bit.

We concluded that the piston corer hit a thick layer of volcanic ash at a shallow depth, which forced the BHA to displace horizontally and bend the core barrel. Because of the shallow penetration (5 mbsf), the sediment provided very little lateral support to the BHA when the the piston corer advanced into the hard layer of ash. The situation can be compared to trying to balance a full bottle of soda in the air with a plastic straw.

Hole 1124C
Because of the incident in Hole 1124B, we decided to drill the first two cores in Hole 1124C with the XCB. Hole 1124C was spudded at 0220 hr on 27 September and washed to 8.0 mbsf where XCB coring was initiated. XCB coring advanced from 8.0 to 27.2 mbsf with 74% recovery. The coring system was then switched to the APC and piston coring advanced from 27.2 to 159.2 mbsf, which was considered APC refusal. The Adara heat-flow shoe was affixed to the APC corer for Cores 181-1124C-5H (55.7 mbsf), 7H (74.7 mbsf), 9H (93.7 mbsf), and 11H (112.7 mbsf). The data from the last three runs were used to compute a temperature gradient of 51.9C/km.

XCB coring deepened the hole from 159.2 to 473.1 mbsf, which was considered the depth objective for this hole.

Logging Operations in Hole 1124C
To prepare for logging, an aluminum go-devil was dropped and the hole swept with 60 barrels of high-viscosity mud. The bit was pulled back in the hole to 447 mbsf and the hole was displaced with 175 barrels of sepiolite mud. The bit was then positioned at the logging depth of 96 mbsf.

Logging operations began at 2100 hr on 29 September and lasted for 19 hr. Logging was conducted from the bottom of the hole at 474 mbsf to the bit at 78 mbsf (picked up from 96 mbsf). Three standard tool-string configurations were run: the triple combination, the FMS-sonic, and the GHMT (Geological High Resolution Magnetic Tool). The NMRS tool on the GHMT failed to operate. The condition of the borehole was good and the quality of the data was excellent. By 1545 hr on 30 September, the Schlumberger logging equipment was disassembled and the drill bit cleared the seafloor at 1630 hr.

Hole 1124D
The vessel was offset 30 m to the north of Hole 1124C. An attempt was made to wash down to 13 mbsf but was given up when very hard formation (possibly cemented ash) was encountered. The vessel was offset an additional 10 m to the west where Hole 1124D was spudded with the XCB at 1917 hr on 30 September. The bit was drilled ahead to 22.6 mbsf with no difficulty and the XCB wash barrel was recovered. APC coring was initiated at this depth and continued to the planned depth objective of 155.6 mbsf.

The drill string was pulled to the surface with the bit clearing the sea floor at 1325 hr. As soon as the bit cleared the seafloor, the beacon was recalled. After the BHA was inspected, the drilling equipment was secured and the vessel was under way to Site 1125, the last site of Leg 181, at 2330 hr on 1 October.



To 181 Summary of Engineering and Drilling Operations: Site 1125

To 181 Table of Contents